OpenStudy (anonymous):

What is the simplified form of 15x^7/9y^5 times 45y^2/5x^3

OpenStudy (tkhunny):

Rearrange using basic properties \[\frac{ 15x^{7} }{ 9y^{5} } \cdot \frac{ 45y^{2} }{ 5x^{3} } = \frac{ 15x^{7} }{ 5x^{3} } \cdot \frac{ 45y^{2} }{ 9y^{5} }\] Does that look any easier?